# Heuristic Methods are Good Teachers to Distill MLPs for Graph Link Prediction

### v1

@@ Appendix A Hyper-parameters
For capped shortest path (CSP) heuristic, we set the upper bound $\tau=6$ for seven smaller datasets and $\tau=2$ for the three larger datasets. When distilling MLPs from teacher models (GNNs or heuristic methods), we follow the same hyperparameter configurations as in [Guo et al., 2023] for sampling context nodes and adopt the configurations from [Li et al., 2024] for training teacher GNNs and non-distilled MLPs. We use 3-layer MLPs for OGBL-Collab, IGB, and OGBL-Citation2, and 2-layer MLPs for all other datasets. Let $\alpha$, $\beta$ denote the loss weights of $L_{R}$ and $L_{D}$, respectively. After distillation, we perform a grid search to determine the loss weights $\alpha$ and $\beta$, where $\alpha,\beta\in\{0,0.001,1,10\}$. Additionally, we use grid search to optimize the margin $\delta$ in $L_{R}$ (Eq. 1), where $\delta\in\{0.05,0.1,0.2\}$. For training ensemble models, we conduct a grid search to find the optimal weight $\lambda$ for L1 regularization (Eq. 7), where $\lambda\in\{0,0.1,1\}$.
Appendix B Additional Experiment Results
B.1 Breakdown of Best GNN and Best LLP
@@
B.2 Additional GNN4LP Teachers
Table 7 shows more performance of GNN4LP models and their distilled MLPs. Specifically, we evaluate NCN [Wang et al., 2024], NCNC [Wang et al., 2024], and BUDDY [Chamberlain et al., 2022]. The results further validate our finding that stronger models might not be better teachers.
@@ B.2 Additional GNN4LP Teachers
Pubmed $47.62\pm 3.22$ $41.17\pm 5.10$ $43.89\pm 5.16$ $46.38\pm 3.09$ $46.73\pm 4.52$ $44.42\pm 3.15$
B.3 Performance of Ensemble Heuristics
Here, we provide more results showing distilling MLPs from ensemble heuristics is sub-optimal. We trained an MLP as a gating function to ensemble different heuristic methods and distilled MLPs from these ensemble heuristics. As shown in Table 8, although ensemble heuristics generally yield higher Hits@K scores, the student MLP distilled from the ensemble heuristics has lower Hits@K. The results suggest that MLPs, limited by their capacity, struggle to effectively learn complex structural information. Consequently, GNNs and GNN4LP models, which integrate various proximities together [Mao et al., 2023, Ma et al., 2024], may also present challenges for MLPs to fully capture.
Table 8: Performance comparison of different heuristic methods and their corresponding MLP-distilled versions across datasets. CC denotes “CN+CSP”, CARC denotes “CC+AA+RA+CSP”. Bolded numbers mark the highest performance. Underlined numbers represent the second highest performance.
Dataset CN CSP CC CARC
@@ B.3 Performance of Ensemble Heuristics
B.4 Complicated Heuristic Teachers
Notably, MLPs cannot learn well from Katz [Katz, 1953] heuristic either. Katz [Katz, 1953] is a well-established heuristic for link prediction, widely effective across various datasets [Li et al., 2024]. Katz is computed as $\sum_{l=1}^{\infty}\lambda^{l}|path^{(l)}(i,j)|$, where $\lambda<1$ is a damping factor and $|path^{(l)}(i,j)|$ is the number of length-$l$ paths between $i$ and $j$. As shown in Table 9, while Katz outperforms CN as a heuristic, the MLP distilled from Katz performs worse than MLPs trained without distillation. It suggests that MLPs cannot learn well from complicated heuristics.
@@ B.4 Complicated Heuristic Teachers
B.5 Effects of the Number of Trainable Parameters
@@
B.6 Problems with “Production” Setting
Guo et al. [Guo et al., 2023] introduced a “production” setting for splitting train/test data in link prediction tasks, where a random subset of nodes is treated as newly added to the graph. These nodes and their associated edges are removed from the training graph. Their results showed a significant performance drop when transitioning from the “transductive” setting to the “production” setting. Similarly, we observe the same performance decline for heuristic-distilled MLPs, to the extent that heuristic distillation may fail to improve MLP performance.
Our analysis reveals that this performance drop is primarily due to the "production" split significantly altering the graph structure, leading to inaccurate heuristic values in the training graph. For instance, triangle counts play a crucial role in link prediction [Adamic and Adar, 2003, Zhang et al., 2021a, Chamberlain et al., 2022]. As shown in Table 11, the number of triangles in the training graphs is drastically reduced under the "production" setting compared to the “transductive” setting. This discrepancy indicates that the removing random nodes make the graphs more “broken” and that “production” setting introduces a substantial structural mismatch between training and test graphs, explaining the performance degradation observed in [Guo et al., 2023].
@@ B.6 Problems with “Production” Setting
Table 11: Triangle counts in training graph under different split settings.
@@
Triangle Counts # Connected Components Triangle Counts # Connected Components
@@
Appendix C Proof of Theorem 3.1
Proof.
By assumption, $s_{i}^{\text{extra}}\perp\!\!\!\perp\bm{x}_{i}$. Thus,
$P(s_{i}^{\text{extra}}\mid\bm{x}_{i})=P(s_{i}^{\text{extra}}).$
So
[…]
Define a GNN teacher:
[…]
This GNN marginalizes over $s^{\text{extra}}$, which is independent of $\bm{x}_{i},\bm{x}_{j}$.
The optimal MLP distilled from $F_{\text{GNN}}$ is:
$\mathbb{E}_{s^{\text{GNN}}}\left[F_{\text{GNN}}\mid\bm{x}_{i},\bm{x}_{j}\right]=\mathbb{E}_{s^{\text{GNN}},s^{\text{extra}}}\left[F_{\text{GNN4LP}}\mid\bm{x}_{i},\bm{x}_{j}\right]=F_{\text{GNN4LP-MLP}}^{*}.$
Thus, the MLP students from $F_{\text{GNN4LP}}$ and $F_{\text{GNN}}$ are identical. ∎

### v2

@@ Appendix F Hyper-parameters
For capped shortest path (CSP) heuristic, we set the upper bound $\tau=6$ for seven smaller datasets and $\tau=2$ for the three larger datasets. When distilling MLPs from teacher models (GNNs or heuristic methods), we follow the same hyperparameter configurations as in [guo2023linkless] for sampling context nodes and adopt the configurations from [heart] for training teacher GNNs and non-distilled MLPs. We use 3-layer MLPs for OGBL-Collab, IGB, and OGBL-Citation2, and 2-layer MLPs for all other datasets. Let $\alpha$, $\beta$ denote the loss weights of $L_{R}$ and $L_{D}$, respectively. After distillation, we perform a grid search to determine the loss weights $\alpha$ and $\beta$, where $\alpha,\beta\in\{0,0.001,1,10\}$. Additionally, we use grid search to optimize the margin $\delta$ in $L_{R}$ (Eq. 8), where $\delta\in\{0.05,0.1,0.2\}$. For training ensemble models, we conduct a grid search to find the optimal weight $\lambda$ for L1 regularization (Eq. 7), where $\lambda\in\{0,0.1,1\}$.
Appendix G Additional Experiment Results
G.1 Inference Speed
One of the primary advantages of our approach is its fast inference training speed. Figure 7 in our paper has shown its advantage in training speed. And prior work, such as LLP, has demonstrated that GNN-to-MLP methods can achieve up to 70 times faster inference compared to GNNs. Our experimental results, as shown in Table 7, confirm that EHDM maintains similar inference speed advantages.
Table 7: Inference time (in ms) for different models on Collab dataset.
Model SAGE SAGE-LLP EHDM
Inference Time (ms) 134.3 2.1 3.6
G.2 Breakdown of Best GNN and Best LLP
@@
G.3 Comparision of GNN4LP and GNN methods
Below we show a more comprehensive comparison between NCN and standard GNNs across seven datasets, including two million-scale graphs. As shown in the Table 10 and Table 11, while NCN consistently achieves the highest accuracy as a teacher, the MLPs distilled from standard GNNs outperform those distilled from NCN in 6 out of 7 cases. Furthermore, our proposed EHDM model consistently achieves the best student performance, while also training 3 times faster.
Table 10: Teacher Model Performance
Dataset GCN SAGE GAT NCN
cora $73.96\pm 1.23$ $71.80\pm 3.08$ $71.77\pm 1.44$ $83.22\pm 1.37$
citeseer $79.91\pm 2.52$ $85.10\pm 2.25$ $80.48\pm 2.21$ $85.45\pm 0.73$
pubmed $68.83\pm 2.84$ $62.45\pm 3.11$ $56.15\pm 3.09$ $72.47\pm 1.86$
CS $64.24\pm 3.52$ $51.20\pm 2.61$ $59.17\pm 4.15$ $74.49\pm 2.37$
computers $27.07\pm 4.45$ $27.07\pm 4.45$ $26.00\pm 3.09$ $39.49\pm 2.30$
coauthor $56.75\pm 1.39$ $9.14\pm 1.64 55$$5.90\pm 1.22$ $63.07$
citation2 $84.90\pm 0.06$ $82.92\pm 0.22$ OOM $89.11$
Table 11: Distilled Model Performance
Dataset GCN-MLP SAGE-MLP GAT-MLP NCN-MLP EHDM
cora $76.62\pm 2.00$ $75.26\pm 3.22$ $74.54\pm 4.34$ $69.37\pm 2.10$ $80.49\pm 1.51$
citeseer $76.53\pm 4.88$ $73.27\pm 4.11$ $74.95\pm 3.14$ $79.29\pm 2.79$ $79.08\pm 1.90$
pubmed $59.95\pm 1.58$ $54.07\pm 4.95$ $58.30\pm 3.66$ $59.85\pm 1.31$ $60.75\pm 3.42$
CS $66.64\pm 1.80$ $65.50\pm 3.91$ $63.65\pm 1.97$ $58.82\pm 4.01$ $74.33\pm 2.59$
computers $23.66\pm 1.61$ $25.80\pm 4.80$ $23.51\pm 2.14$ $23.01\pm 6.21$ $30.41\pm 2.90$
collab $49.30\pm 0.79$ $48.00\pm 0.62$ $48.45\pm 0.82$ $48.00\pm 1.79$ $49.27\pm 0.88$
citation2 $42.78\pm 0.10$ $42.62\pm 0.08$ NA $42.58\pm 2.10$ $46.58\pm 0.10$
G.4 Additional GNN4LP Teachers
Table 12 shows more performance of GNN4LP models and their distilled MLPs. Specifically, we evaluate NCN [ncnc], NCNC [ncnc], and BUDDY [buddy]. The results further validate our finding that stronger models might not be better teachers.
@@ G.4 Additional GNN4LP Teachers
Pubmed $47.62\pm 3.22$ $41.17\pm 5.10$ $43.89\pm 5.16$ $46.38\pm 3.09$ $6.73\pm 4.52 44$$4.42\pm 3.15$
G.5 Performance of Ensemble Heuristics
Here, we provide more results showing distilling MLPs from ensemble heuristics is sub-optimal. We trained an MLP as a gating function to ensemble different heuristic methods and distilled MLPs from these ensemble heuristics. As shown in Table 13, although ensemble heuristics generally yield higher Hits@K scores, the student MLP distilled from the ensemble heuristics has lower Hits@K. The results suggest that MLPs, limited by their capacity, struggle to effectively learn complex structural information. Consequently, GNNs and GNN4LP models, which integrate various proximities together [mao2023revisiting, linkmoe], may also present challenges for MLPs to fully capture.
Table 13: Performance comparison of different heuristic methods and their corresponding MLP-distilled versions across datasets. $C$ denotes “CN+CSP”, CARC denotes “CC+AA+RA+CSP”. Bolded numbers mark the highest performance. Underlined numbers represent the second highest performance.
Dataset CN CSP $C$ CARC
@@ G.5 Performance of Ensemble Heuristics
G.6 Complicated Heuristic Teachers
Notably, MLPs cannot learn well from Katz [katz1953new] heuristic either. Katz [katz1953new] is a well-established heuristic for link prediction, widely effective across various datasets [heart]. Katz is computed as $\sum_{l=1}^{\infty}\lambda^{l}|path^{(l)}(i,j)|$, where $\lambda<1$ is a damping factor and $|path^{(l)}(i,j)|$ is the number of length-$l$ paths between $i$ and $j$. As shown in Table 14, while Katz outperforms CN as a heuristic, the MLP distilled from Katz performs worse than MLPs trained without distillation. It suggests that MLPs cannot learn well from complicated heuristics.
@@ G.6 Complicated Heuristic Teachers
G.7 Effects of the Number of Trainable Parameters
@@
G.8 Problems with “Production” Setting
Guo et al. [guo2023linkless] introduced a “production” setting for splitting train/test data in link prediction tasks, where a random subset of nodes is treated as newly added to the graph. These nodes and their associated edges are removed from the training graph. Their results showed a significant performance drop when transitioning from the “transductive” setting to the “production” setting. Similarly, we observe the same performance decline for heuristic-distilled MLPs, to the extent that heuristic distillation may fail to improve MLP performance.
Our analysis reveals that this performance drop is primarily due to the "production" split significantly altering the graph structure, leading to inaccurate heuristic values in the training graph. For instance, triangle counts play a crucial role in link prediction [adamic2003friends, zhang2021labeling, buddy]. As shown in Table 16, the number of triangles in the training graphs is drastically reduced under the "production" setting compared to the “transductive” setting. This discrepancy indicates that the removing random nodes make the graphs more “broken” and that “production” setting introduces a substantial structural mismatch between training and test graphs, explaining the performance degradation observed in [guo2023linkless].
@@ G.8 Problems with “Production” Setting
Table 16: Triangle counts and the number of connected components (#$C$) in training graph under different split settings.
@@
Triangle Counts #$C$ Triangle Counts #$C$
@@
Appendix H Limitations
Since MLPs rely heavily on rich node features to perform effectively, our experiments focus primarily on datasets from social networks and recommendation systems, where such features are abundant. In contrast, biological graphs, such as protein-protein interaction networks, are not included in our evaluation, as they typically lack informative node features, making them less suitable for MLP-based approaches. This limitation is not specific to our method but is inherent to all MLP distillation methods on graphs.